# Uncertainty-Aware Hybrid Inference with On-Device Small and Remote Large Language Models

### v2

@@ III-A Feasibility of Predicting Rejection Probability Using Uncertainty
Our approach is inspired by the concept of uncertainty, which quantifies a model’s confidence in its output. Specifically, we hypothesize that by measuring the SLM’s uncertainty regarding a generated draft token, we can predict the rejection probability of that token in the LLM without requiring the target probability. We conduct the following experiment to explore the feasibility of the above hypothesis. Before going further, we denote the rejection probability and uncertainty of the $t$-th round of HLM inference as $\beta(t)\coloneqq\max\left(1-\frac{y_{d}(t)}{x_{d}(t)},0\right)$ and $u(t)$, respectively.
@@
$\beta(t)=a\cdot u(t)+b,\quad\forall t.$ (5)
@@
III-B Uncertainty-Aware Hybrid Language Model
@@ III-C Guideline for Setting Uncertainty Threshold
This subsection outlines the design of an uncertainty threshold for selectively skipping uplink transmissions. Specifically, tokens are skipped if their rejection probability $\beta(t)$ satisfies $\beta(t)\leq\beta_{\text{th}}$, where $\beta_{\text{th}}$ denotes the rejection probability threshold. Two skipping strategies are considered: risk-averse skipping, which skips only tokens predicted to be immediately accepted ($\beta_{\text{th}}=0$), and risk-prone skipping, which skips both tokens predicted to be immediately accepted and those probabilistically accepted. These strategies involve a trade-off between inference accuracy and token throughput. Risk-averse skipping is lossless, preserving a vocabulary distribution consistent with LLM inference, as expressed by:
$(1-\beta(t))\cdot x_{v}(t)+\beta(t)\cdot P_{v}(t)=y_{v}(t),\quad\forall v,t,$ (9)
with the proof of this equality provided in [7]. However, this approach does not facilitate higher token throughput. In contrast, risk-prone skipping increases token throughput by deliberately relaxing the equality in (9), introducing a bias $\beta_{\text{th}}>0$ to $\beta(t)$. This adjustment, however, may come at the expense of reduced inference accuracy. In this context, this subsection focuses on risk-prone skipping, aiming to maximize token throughput, derive its uncertainty threshold, and examine the associated risks.
Revisiting the rejection scenario in U-HLM, rejection occurs when the target probability $y_{d}(t)$ is less than the draft probability $x_{d}(t)$, with a probability of $1-y_{d}(t)/x_{d}(t)$. For the U-HLM device aiming to skip uplink transmissions, $x_{d}(t)$ is deterministic on the device, while $y_{d}(t)$ is treated as a random variable. Under these conditions, the rejection probability $\beta(t)$ can be reformulated as follows, incorporating the stochastic nature of $y_{d}(t)$:
$\beta(t)=P(y_{d}(t)<x_{d}(t))\cdot\textsf{E}_{y_{d}(t)}\left[\frac{y_{d}(t)}{x_{d}(t)}\mid y_{d}(t)<x_{d}(t)\right].$ (10)
Ideally, the uncertainty threshold should be tailored to each token individually. However, for simplicity, we use a fixed threshold. To determine this threshold, we derive a bound on the uncertainty under the assumption that the tokens are independently and identically distributed (i.i.d.) as follows:
@@
Assuming that both the uncertainty and the rejection probability are i.i.d. across rounds, with $u\coloneqq u(t)$ and $\beta\coloneqq\beta(t)$ for any $t$, the uncertainty threshold $u_{\text{th}}$ in U-HLM is derived as:
$u_{\text{th}}=\frac{\Delta-b}{a}.$ (11)
@@
which completes the proof of (11). The upper bound of $R$ is obtained by applying Hölder’s inequality. ∎
In (14), applying the threshold at the upper bound corresponds to risk-prone skipping, while applying it at the lower bound corresponds to risk-averse skipping. In Theorem 1, $R$ represents the expected increase in rejection probability when adopting risk-prone skipping instead of risk-averse skipping, reflecting the density of uncertainty. This captures the theoretical impact on U-HLM’s inference accuracy, which is practically affected when an uplink for a token at risk of rejection is skipped, potentially resulting in a loss of accuracy. The impact on actual inference accuracy, along with its theoretical upper bound derived in (13), is thoroughly validated through the experiments presented in Section IV.
@@ IV-A Validation for Theorem 1
Next, we derive the upper bound of $R$. Instead of using the continuous PDF, we utilize the discrete empirical probability density of uncertainty shown in Figure 5 (left), yielding $R<4\times 10^{-3}$. Experimentally, we observe $R=2.24\times 10^{-3}$, confirming that the upper bound is satisfied. This result aligns with the minimal cosine similarity loss observed at the uncertainty threshold for risk-prone skipping, as shown in Figure 4, which can be attributed to the low probability values of uncertainty between the two thresholds, as illustrated in Figure 5. Based on these findings, we set $u_{\text{th}}=0.431$.
@@ IV-B Inference Accuracy and Token Throughput Comparisons
Table I and Figure 6 compare the inference accuracy and token throughput of U-HLM with other inference methods. Table I demonstrates that U-HLM outperforms both SLM and Rand-HLM in terms of cosine similarity across all datasets. For instance, with the Alpaca dataset, while Rand-HLM improves cosine similarity by only 2.08% over SLM, U-HLM achieves a 5.08% improvement—representing a 244% greater accuracy gain. On average, U-HLM achieves up to 97.54% and 100.05% of the inference accuracy of LLM and HLM, respectively. As shown in Figure 6 (right), U-HLM delivers the highest token throughput across all SNRs, second only to SLM, with a particularly notable 2.54$\times$ increase in token throughput compared to HLM without skipping. This advantage stems from U-HLM’s ability to accurately and frequently skip uplinks for tokens while maintaining inference quality. This is further evidenced by its higher TSR and lower TR compared to Rand-HLM in Figure 6 (left), reflecting a 45.93% reduction in uplink transmissions and LLM computations. Moreover, U-HLM demonstrates substantial improvements in TSR, TR, and token throughput, particularly when temperature perturbation is applied and under low average SNR conditions.
@@ Acknowledgments
This work was supported in part by the Korean Ministry of Science and ICT (MSIT) under the National Research Foundation of Korea (NRF) grant (No. 2023-11-1836), in part by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by MSIT (No. RS-2024-00404972), and in part by IITP under YKCS Open RAN Global Collaboration Center (IITP-2024-RS-2024-00434743) grant funded by MSIT.

### v3

@@ III-A Feasibility of Predicting Rejection Probability Using Uncertainty
Our approach is inspired by the concept of uncertainty, which quantifies a model’s confidence in its output. Specifically, we hypothesize that by measuring the SLM’s uncertainty regarding a generated draft token, we can predict the rejection probability of that token in the LLM without requiring the target probability. We conduct the following experiment to explore the feasibility of the above hypothesis. Before proceeding, we define the rejection probability and uncertainty for the draft token $d$ in the $t$-th round of HLM inference as $\beta_{d}(t)\coloneqq\max\left(1-\frac{y_{d}(t)}{x_{d}(t)},0\right)$ and $u(t)$, respectively.
@@
$\beta_{d}(t)=a\cdot u(t)+b,\quad\forall t.$ (5)
@@
III-B Uncertainty-Aware Opportunistic Hybrid Language Model
@@ III-C Guideline for Setting Uncertainty Threshold
This subsection outlines the design of an uncertainty threshold for selectively skipping uplink transmissions. Specifically, the device aims to omit uplink transmission for draft tokens whose predicted rejection probability $\beta_{d}(t)$ meets the condition $\beta_{d}(t)\leq\beta_{\text{th}}$, where $\beta_{\text{th}}$ represents the rejection probability threshold. Two skipping strategies are considered: risk-averse skipping, which skips only tokens predicted to be immediately accepted ($\beta_{\text{th}}=0$), and risk-prone skipping, which skips both tokens predicted to be immediately accepted and those probabilistically accepted. These strategies involve a trade-off between inference accuracy and token throughput. Risk-averse skipping is lossless, ensuring that the vocabulary distribution remains consistent with LLM inference. To formalize this, we extend the definition of rejection probability to any vocabulary token $v$ at time $t$ as: $\beta_{v}(t)\coloneqq\max\left(1-\frac{y_{v}(t)}{x_{v}(t)},0\right)$. Using this, the preserved vocabulary distribution under risk-averse skipping satisfies the unbiasedness requirement, expressed as:
$x_{v}(t)\cdot(1-\beta_{v}(t))+\sum_{i=1}^{|\mathcal{V}|}{(x_{i}(t)\cdot\beta_{i}(t))}\cdot P_{v}(t)=y_{v}(t),\quad\forall v,t,$ (9)
with the proof of this equality provided in [7]. However, this approach does not facilitate higher token throughput. In contrast, risk-prone skipping increases token throughput by deliberately relaxing the equality in (9), introducing a bias $\beta_{\text{th}}>0$ to $\beta_{v}(t)$. This adjustment, however, may come at the expense of reduced inference accuracy. In this context, this subsection focuses on risk-prone skipping, aiming to maximize token throughput, derive its uncertainty threshold, and examine the associated risks.
Revisiting the rejection scenario in U-HLM, rejection occurs when the target probability $y_{d}(t)$ is less than the draft probability $x_{d}(t)$, with a probability of $1-y_{d}(t)/x_{d}(t)$. For the U-HLM device aiming to skip uplink transmissions, $x_{d}(t)$ is deterministic on the device, while $y_{d}(t)$ is treated as a random variable. Under these conditions, the rejection probability $\beta_{d}(t)$ can be reformulated as follows, incorporating the stochastic nature of $y_{d}(t)$:
$\beta_{d}(t)=P(y_{d}(t)<x_{d}(t))\cdot\textsf{E}_{y_{d}(t)}\left[\frac{y_{d}(t)}{x_{d}(t)}\mid y_{d}(t)<x_{d}(t)\right].$ (10)
Ideally, the uncertainty threshold should be adjusted individually for each $t$-th round. However, for simplicity, we adopt a fixed threshold. To determine this threshold, we derive an uncertainty bound under the assumption that uncertainty and rejection probability are independently and identically distributed (i.i.d.) across rounds:
@@
Under the i.i.d. assumption, with $u\coloneqq u(t)$ and $\beta\coloneqq\beta(t)$ for any $t$, the uncertainty threshold $u_{\text{th}}$ in U-HLM is given by:
$u_{\text{th}}=\frac{\Delta-b}{a},$ (11)
where $\Delta=P(y_{d}<x_{d})$ represents the probability that a draft token $d$ is either probabilistically accepted or rejected.
@@
which concludes the derivation of (11). The upper bound of $R$ is obtained by applying Hölder’s inequality. ∎
In (15), setting the threshold at the upper bound corresponds to risk-prone skipping, while setting it at the lower bound corresponds to risk-averse skipping. In Theorem 1, $R$ quantifies the expected increase in rejection probability when adopting risk-prone skipping instead of risk-averse skipping, reflecting the density of uncertainty. This captures the theoretical impact on U-HLM’s inference accuracy, where its loss arises when skipping the uplink for tokens at risk of rejection. This impact, along with the expected rejection risk and its upper bound derived in Theorem 1, is thoroughly validated through the experiments presented in Section IV.
@@ IV-A Validation for Theorem 1
Next, we derive the upper bound of $R$. Instead of using the continuous PDF, we utilize the discrete empirical probability density of uncertainty shown in Figure 5 (left), yielding $R<4\times 10^{-3}$ from (14). Experimentally, we observe $R=2.24\times 10^{-3}$, confirming that the upper bound is satisfied. This small expected rejection risk $R$ aligns with the minimal cosine similarity loss observed at the uncertainty threshold for risk-prone skipping, as shown in Figure 4. This minimal loss can be attributed to the low probability values of uncertainty between the two thresholds, as illustrated in Figure 5. Based on these findings, we adopt an uncertainty threshold of $u_{\text{th}}=0.431$ for risk-prone skipping.
@@ IV-B Inference Accuracy and Token Throughput Comparisons
Table I and Figure 6 compare the inference accuracy and token throughput of U-HLM with other inference methods. Table I demonstrates that U-HLM outperforms both SLM and Rand-HLM in terms of cosine similarity across all datasets. For instance, with the Alpaca dataset, while Rand-HLM improves cosine similarity by only 2.08% over SLM, U-HLM achieves a 5.08% improvement—representing a 244% greater accuracy gain. On average, U-HLM achieves up to 97.54% and 100.05% of the inference accuracy of LLM and HLM, respectively. As shown in Figure 6 (right), U-HLM delivers the highest token throughput across all SNRs, second only to SLM, with a particularly notable 2.54$\times$ increase in token throughput compared to HLM without skipping.
This advantage stems from U-HLM’s ability to accurately and frequently skip uplinks for tokens while maintaining inference quality. This is further evidenced by its higher TSR and lower TR compared to Rand-HLM in Figure 6 (left), reflecting a 45.93% reduction in uplink transmissions and LLM computations. Moreover, U-HLM demonstrates substantial improvements in TSR, TR, and token throughput, particularly when temperature perturbation is applied and under low average SNR conditions. This is because, as shown in Figure 3, temperature perturbation facilitates accurate rejection probability estimation due to its linearity, which exhibits a higher slope over a broader uncertainty range. Notably, this increased slope also reduces the upper bound of $R$ in (14), potentially enhancing inference accuracy.
@@ Acknowledgments
This work was supported in part by the Korean Ministry of Science and ICT (MSIT) under the National Research Foundation of Korea (NRF) grant (No. 2023-11-1836), in part by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by MSIT (No. RS-2024-00404972), in part by IITP under YKCS Open RAN Global Collaboration Center (IITP-2024-RS-2024-00434743) grant funded by MSIT, in part by IITP-ITRC (Information Technology Research Center) grant funded by MSIT (IITP-2025-RS-2023-00259991), and in part by SUTD Kickstarter Initiative (SKI 2021_06_08).